Please help me verify this identity, im stuck!
sec x + csc (-x) / sec x csc x = sin x - cos x
thanks!

Answer by stanbon(75887)

sec x + csc (-x) / sec x csc x = sin x - cos x
----
Cross-multiply to get:
sec(x) + csc(-x) = (sec(x)*csc(x))(sin(x)-cos(x))
------
Distribute on the right side and,
Since csc(-x) = -csc(x) on the left side you get::
sec(x) - csc(x) = sec(x) - csc(x)
===========
Cheers,
Stan H.
===========